Fill in the blank with the term that would complete the square.

a)x^2-8x+
b)x^2-7x+
c)x^2-2/3x+
d)x^2-px+

Answer by norymae(15)

formula:
ax^2+bx+c=0
divide b by 2 then square the answer:
a. 8/2=4^2=16
b.7/2=(7/2)^2=49/4
c.(2/3)/2=(1/3)^2=1/9
d.p/2=(p/2)^2=p^2/4